# Conformal Defects in Neural Network Field Theories

### v1

@@ IV.2 Reciprocal NN-FTs
$\mathbb{E}[\Phi_{\Delta_{1}}(X_{1})\Phi_{\Delta_{2}}(X_{2})]$ $=\frac{2^{\frac{\Delta_{2}-\Delta_{1}}{2}}\sec(\pi\Delta_{1})}{\Gamma(\frac{\Delta_{1}-\Delta_{2}}{2}+1)\Gamma(\Delta_{2})}(\hat{\sigma}^{2}X_{1}\bullet X_{2}+\tilde{\sigma}^{2}X_{1}\circ X_{2})^{-\Delta_{1}}\tilde{\sigma}^{\Delta_{1}-\Delta_{2}}\|X_{2}\|_{\circ}^{\Delta_{1}-\Delta_{2}}\times$
$\qquad\qquad\qquad\times{}_{2}F_{1}\left(\frac{\Delta_{1}}{2},\frac{\Delta_{1}+1}{2};\frac{\Delta_{1}-\Delta_{2}}{2}+1;\frac{\tilde{\sigma}^{4}\|X_{1}\|_{\circ}^{2}\|X_{2}\|_{\circ}^{2}}{\left(\tilde{\sigma}^{2}X_{1}\circ X_{2}+\hat{\sigma}^{2}X_{1}\bullet X_{2}\right)^{2}}\right),$ (59)
@@
$\mathbb{E}[\Phi_{\Delta_{1}}(X_{1})\Phi_{\Delta_{2}}(X_{2})]$ $=\frac{2^{\frac{\Delta_{2}-\Delta_{1}}{2}}\sec(\pi\Delta_{1})\tilde{\sigma}^{-\Delta_{1}-\Delta_{2}}}{\Gamma(\frac{\Delta_{1}-\Delta_{2}}{2}+1)\Gamma(\Delta_{2})}\frac{(-\frac{1}{2}\frac{\hat{\sigma}^{2}}{\tilde{\sigma}^{2}}\chi+\cos(\psi))^{-\Delta_{1}}}{\|X_{1}\|_{\circ}^{\Delta_{1}}\|X_{2}\|_{\circ}^{\Delta_{2}}}\times$
$\qquad\times{}_{2}F_{1}\left(\frac{\Delta_{1}}{2},\frac{\Delta_{1}+1}{2};\frac{\Delta_{1}-\Delta_{2}}{2}+1;\left(-\frac{1}{2}\frac{\hat{\sigma}^{2}}{\tilde{\sigma}^{2}}\chi+\cos(\psi)\right)^{-2}\right)~.$ (60)
@@ IV.3 Positivity in reciprocal NN-FTs
with $\rho=|\Theta|$, $\check{\Theta}\in\mathds{S}^{D-1}$, and $f(\rho)\geq 0$, and where $w(\check{\Theta})\geq 0$ is normalized to unity with the assumption that $p$ is smooth and strictly positive in a neighbourhood of $\Theta=0$.
@@ V Conclusion
  • Global symmetries and monodromy defects. As briefly touched upon in § II, implementing global symmetries in a field-theoretic way in NN-FTs requires a novel approach. That is, we would naturally draw the parallel from global symmetry in field space affected by acting on a collection of fields with some group action where the action functional and measure in the path integral, if the symmetry is non-anomalous, remain invariant to a NN-FT where the group acts on an ensemble of networks. In order for the NN-FT to be invariant under the symmetry, we would need to require that the generalized measure in the ensemble $\int D\Theta P(\Theta)$ and the expectations computed with respect to it remain invariant.
@@
  • Two-dimensional theories. Given the flexibility of conformal fields in NN-FTs in that they can be defined in arbitrary dimensions, it is natural to ask how to construct novel two-dimensional conformal fields and whether they can realize Virasoro symmetry. We will address this problem in upcoming work Capuozzo:2025xxx ; following in future work with an eye toward extending our defect construction to explore boundary physics in two-dimensions, e.g. Ishibashi states, from the perspective of NNs.
@@ Appendix A Useful identities
When $\lambda=1$, we explicitly find $\mathbb{E}[(X\bullet\theta)^{2}]=X\bullet X\,\hat{\sigma}^{2}$. Let us now assume that $\mathbb{E}[(X\bullet\theta)^{2\lambda}]=\frac{(2\lambda-1)!}{2^{\lambda-1}(\lambda-1)!}(X\bullet X)^{\lambda}\hat{\sigma}^{2\lambda}$ for some $\lambda\in\mathbb{N}^{*}$. Then, using Stein’s lemma (with $\hat{\mu}_{1}=0$ and $\hat{\Sigma}_{AB}=\hat{\sigma}^{2}\delta_{AB}$) ,
@@
$=(2\lambda+1)(X\bullet X)\hat{\sigma}^{2}\mathbb{E}[(X\bullet X)^{2\lambda}]$
$=\frac{(2\lambda+1)!}{2^{\lambda}\lambda!}(X\bullet X)^{\lambda+1}\hat{\sigma}^{2(\lambda+1)}~.$ (83)
@@
$\mathbb{E}[(X_{1}\bullet\theta)^{n_{1}}(X_{2}\bullet\theta)^{n_{2}}]$ $=\sum_{\begin{subarray}{c}n=0\\ n_{1}-n=0\,[2]\\ n_{2}-n=0\,[2]\end{subarray}}^{\operatorname{min}(n_{1},n_{2})}(X_{1}\bullet X_{2})^{n}(X_{1}\bullet X_{1})^{\frac{n_{1}-n}{2}}(X_{2}\bullet X_{2})^{\frac{n_{2}-n}{2}}\hat{\sigma}^{n_{1}+n_{2}}\frac{2^{\frac{2n-n_{1}-n_{2}}{2}}\Gamma(n_{1}+1)\Gamma(n_{2}+1)}{\Gamma(n+1)\Gamma(\frac{n_{1}-n+2}{2})\Gamma(\frac{n_{2}-n+2}{2})}$ (84)
for all $n_{1},n_{2}\in\mathbb{N}$. A similar expression holds for the tangent space random variables $X_{1}\circ\theta$, where $\bullet$ should be replaced by $\circ$ and $\tilde{\sigma}^{2}$ by $\hat{\sigma}^{2}$.
@@
$\mathbb{E}[(X_{1}\bullet\theta)^{n_{1}}(X_{2}\bullet\theta)^{n_{2}}]$ $=(X_{1}\bullet X_{1})^{\frac{n_{1}}{2}}(X_{2}\bullet X_{2})^{\frac{n_{2}}{2}}\left(\frac{\hat{\sigma}^{2}}{2}\right)^{\frac{n_{1}+n_{2}}{2}}\frac{\Gamma(n_{1}+1)\Gamma(n_{2}+1)}{\Gamma\left(\frac{n_{1}}{2}+1\right)\Gamma\left(\frac{n_{2}}{2}+1\right)}\times$
$\qquad\qquad\times{}_{2}F_{1}\left(-\frac{n_{1}}{2},-\frac{n_{2}}{2};\frac{1}{2};\frac{(X_{1}\bullet X_{2})^{2}}{X_{1}\bullet X_{1}X_{2}\bullet X_{2}}\right)~,$ (86a)
@@
$\mathbb{E}[(X_{1}\bullet\theta)^{n_{1}}(X_{2}\bullet\theta)^{n_{2}}]$ $=(X_{1}\bullet X_{2})(X_{1}\bullet X_{1})^{\frac{n_{1}-1}{2}}(X_{2}\bullet X_{2})^{\frac{n_{2}-1}{2}}2\left(\frac{\hat{\sigma}^{2}}{2}\right)^{\frac{n_{1}+n_{2}}{2}}\frac{\Gamma(n_{1}+1)\Gamma(n_{2}+1)}{\Gamma\left(\frac{n_{1}-1}{2}+1\right)\Gamma\left(\frac{n_{2}-1}{2}+1\right)}\times$
$\qquad\qquad\times{}_{2}F_{1}\left(-\frac{n_{1}-1}{2},-\frac{n_{2}-1}{2};\frac{3}{2};\frac{(X_{1}\bullet X_{2})^{2}}{X_{1}\bullet X_{1}X_{2}\bullet X_{2}}\right)~,$ (86b)

### v2

@@ 4.2 Reciprocal NN-FTs
$\mathbb{E}[\Phi_{\Delta_{1}}(X_{1})\Phi_{\Delta_{2}}(X_{2})]$ $=\frac{2^{\frac{\Delta_{2}-\Delta_{1}}{2}}\sec(\pi\Delta_{1})}{\Gamma(\frac{\Delta_{1}-\Delta_{2}}{2}+1)\Gamma(\Delta_{2})}\frac{(\hat{\sigma}^{2}X_{1}\bullet X_{2}+\tilde{\sigma}^{2}X_{1}\circ X_{2})^{-\Delta_{1}}}{\tilde{\sigma}^{\Delta_{2}-\Delta_{1}}\|X_{2}\|_{\circ}^{\Delta_{2}-\Delta_{1}}}\times$ (59)
$\qquad\times{}_{2}F_{1}\left(\frac{\Delta_{1}}{2},\frac{\Delta_{1}+1}{2};\frac{\Delta_{1}-\Delta_{2}}{2}+1;\frac{\tilde{\sigma}^{4}\|X_{1}\|_{\circ}^{2}\|X_{2}\|_{\circ}^{2}}{\left(\tilde{\sigma}^{2}X_{1}\circ X_{2}+\hat{\sigma}^{2}X_{1}\bullet X_{2}\right)^{2}}\right),$
@@
$\mathbb{E}[\Phi_{\Delta_{1}}(X_{1})\Phi_{\Delta_{2}}(X_{2})]$ $=\frac{2^{\frac{\Delta_{2}-\Delta_{1}}{2}}\sec(\pi\Delta_{1})\tilde{\sigma}^{-\Delta_{1}-\Delta_{2}}}{\Gamma(\frac{\Delta_{1}-\Delta_{2}}{2}+1)\Gamma(\Delta_{2})}\frac{(-\frac{1}{2}\frac{\hat{\sigma}^{2}}{\tilde{\sigma}^{2}}\chi+\cos(\psi))^{-\Delta_{1}}}{\|X_{1}\|_{\circ}^{\Delta_{1}}\|X_{2}\|_{\circ}^{\Delta_{2}}}\times$ (60)
$\qquad\times{}_{2}F_{1}\left(\frac{\Delta_{1}}{2},\frac{\Delta_{1}+1}{2};\frac{\Delta_{1}-\Delta_{2}}{2}+1;\left(-\frac{1}{2}\frac{\hat{\sigma}^{2}}{\tilde{\sigma}^{2}}\chi+\cos(\psi)\right)^{-2}\right)~.$
@@ 4.3 Positivity in reciprocal NN-FTs
with $\rho=|\Theta|$, $\check{\Theta}\in\mathds{S}^{D-1}$, and $f(\rho)\geq 0$, and where $w(\check{\Theta})\geq 0$ is normalized to unity with the assumption that $p$ is smooth and strictly positive in a neighborhood of $\Theta=0$.
@@ 5 Conclusion
  • Global symmetries and monodromy defects. As briefly touched upon in § 2, implementing global symmetries in a field-theoretic way in NN-FTs requires a novel approach. That is, we would naturally draw the parallel from global symmetry in field space implemented by acting on a collection of fields with some group action where the action functional and measure in the path integral, if the symmetry is non-anomalous, remain invariant to a NN-FT where the group acts on an ensemble of networks. In order for the NN-FT to be invariant under the symmetry, we would need to require that the generalized measure in the ensemble $\int D\Theta P(\Theta)$ and the expectations computed with respect to it remain invariant.
@@
  • Two-dimensional theories. Given the flexibility of conformal fields in NN-FTs in that they can be defined in arbitrary dimensions, it is natural to ask how to construct novel two-dimensional conformal fields and whether they can realize Virasoro symmetry. We will address this problem in upcoming work Robinson:2025ybg ; with an eye toward extending our defect construction to explore boundary physics in two-dimensions, e.g. Ishibashi states, from the perspective of NNs.
@@ Appendix A Useful identities
When $\lambda=1$, we explicitly find $\mathbb{E}[(X\bullet\theta)^{2}]=\|X\|_{\bullet}^{2}\,\hat{\sigma}^{2}$. Let us now assume that $\mathbb{E}[(X\bullet\theta)^{2\lambda}]=\frac{(2\lambda-1)!}{2^{\lambda-1}(\lambda-1)!}\|X\|_{\bullet}^{2\lambda}\hat{\sigma}^{2\lambda}$ for some $\lambda\in\mathbb{N}^{*}$. Then, using Stein’s lemma (with $\hat{\mu}_{1}=0$ and $\hat{\Sigma}_{AB}=\hat{\sigma}^{2}\delta_{AB}$) ,
@@
$=(2\lambda+1)\|X\|_{\bullet}^{2}\hat{\sigma}^{2}\mathbb{E}[\|X\|_{\bullet}^{4\lambda}]$
$=\frac{(2\lambda+1)!}{2^{\lambda}\lambda!}\|X\|_{\bullet}^{2\lambda+2}\hat{\sigma}^{2(\lambda+1)}~.$ (83)
@@
$\mathbb{E}[(X_{1}\bullet\theta)^{n_{1}}(X_{2}\bullet\theta)^{n_{2}}]=\sum_{\begin{subarray}{c}n=0\\ n_{1}-n=0\,[2]\\ n_{2}-n=0\,[2]\end{subarray}}^{\operatorname{min}(n_{1},n_{2})}$ $(X_{1}\bullet X_{2})^{n}\|X_{1}\|_{\bullet}^{n_{1}-n}\|X_{2}\|_{\bullet}^{n_{2}-n}\hat{\sigma}^{n_{1}+n_{2}}\times$ (84)
$\hskip 7.11317pt\times\frac{2^{\frac{2n-n_{1}-n_{2}}{2}}\Gamma(n_{1}+1)\Gamma(n_{2}+1)}{\Gamma(n+1)\Gamma(\frac{n_{1}-n+2}{2})\Gamma(\frac{n_{2}-n+2}{2})}$
for all $n_{1},n_{2}\in\mathbb{N}$. A similar expression holds for the tangent space random variables $X_{1}\circ\theta$, where $\bullet$ should be replaced by $\circ$ and $\hat{\sigma}^{2}$ by $\tilde{\sigma}^{2}$.
@@
$\mathbb{E}[(X_{1}\bullet\theta)^{n_{1}}(X_{2}\bullet\theta)^{n_{2}}]$ $=\|X_{1}\|_{\bullet}^{n_{1}}\|X_{2}\|_{\bullet}^{n_{2}}\left(\frac{\hat{\sigma}^{2}}{2}\right)^{\frac{n_{1}+n_{2}}{2}}\frac{\Gamma(n_{1}+1)\Gamma(n_{2}+1)}{\Gamma\left(\frac{n_{1}}{2}+1\right)\Gamma\left(\frac{n_{2}}{2}+1\right)}\times$
$\qquad\times{}_{2}F_{1}\left(-\frac{n_{1}}{2},-\frac{n_{2}}{2};\frac{1}{2};\frac{(X_{1}\bullet X_{2})^{2}}{\|X_{1}\|_{\bullet}^{2}\|X_{2}\|_{\bullet}^{2}}\right)~,$ (86a)
@@
$\mathbb{E}[(X_{1}\bullet\theta)^{n_{1}}(X_{2}\bullet\theta)^{n_{2}}]$ $=\frac{X_{1}\bullet X_{2}}{\|X_{1}\|_{\bullet}^{1-n_{1}}\|X_{2}\|_{\bullet}^{1-n_{2}}}\hat{\sigma}^{n_{1}+n_{2}}\frac{2^{\frac{2-n_{1}-n_{2}}{2}}\Gamma(n_{1}+1)\Gamma(n_{2}+1)}{\Gamma\left(\frac{n_{1}-1}{2}+1\right)\Gamma\left(\frac{n_{2}-1}{2}+1\right)}\times$ (86b)
$\qquad\times{}_{2}F_{1}\left(\frac{1-n_{1}}{2},\frac{1-n_{2}}{2};\frac{3}{2};\frac{(X_{1}\bullet X_{2})^{2}}{\|X_{1}\|_{\bullet}^{2}\|X_{2}\|_{\bullet}^{2}}\right)~,$